OpenStudy (anonymous):

Factor the Polynomials completely: 4x^2+5x-6

OpenStudy (anonymous):

x=(-5+-sqrt(25-4(4)(-6)))/(2(4))=(-5+-11)/8 x=-2 and x=3/4

OpenStudy (anonymous):

i used the quadratic formula for this question

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

welcome